OpenStudy (anonymous):

the numerator of a fraction is 5 less than the denominator .if 1 is added to both numerator and denominator,the fraction would become 2\3 .find the fraction.

OpenStudy (anonymous):

let denominator be x. given fraction is x-5/x adding 1 to both, x-4/x+1 = 2/3 2x+2 = 3x-12 x=14

OpenStudy (anonymous):

3/1 right?